### v1

@@ 4 Conclusion and Outlook
We have investigated the impact of different types of measurement uncertainties on the inferred stability of artificial, intrinsically stable planes of satellite galaxies. Our key results are:
@@
  • At the level of minimum Gaia uncertainties (with $\epsilon_{\mu}$ = 0.04 mas yr-1, similar to Gaia systematics), the inferred plane widths begin to diverge from true one towards larger values. While the thin initial plane remains relatively well constrained, its inferred width relative to the true plane almost doubles within 3 Gyr. As we increase the applied proper motion uncertainties ( $\epsilon_{\mu}$ = 0.08 and 0.12 mas yr-1), the inferred plane widths increase substantially, and one might infer that the intrinsically stable satellite planes are highly unstable. The $c/a$ shows a linear increase with the addition of more uncertainties.
@@
  • When performing backward integration with three different intrinsic plane height ($h=10,20$, and 30 kpc), we find that applying realistic uncertainties in the proper motion introduce a substantial broadening in the plane of satellites in all cases. The apparent broadening effect we identified is therefore not sensitive to the exact intrinsic plane height assumed for the plane.
@@
  • When applying 5% distance uncertainties, a noticeable effect is observed which increases the plane width during backward-integration relative to the intrinsic planes’ width. However, this does not dominate the inferred evolution of the satellite plane. Additionally, the impact of distance uncertainties becomes less important with higher levels of proper motion uncertainties.
@@
  • The influence of adopting incorrect Milky Way potential models leads to an apparent widening of the inferred past satellite planes. The impact is especially significant for $MW_{\text{low}}$, which has a 40% smaller halo mass compared to the fiducial model considered for our intrinsic satellite planes. Due to its lower mass, it struggles to maintain cohesion among the test satellites, leading to increased radial distances and an extension of its major, minor, and intermediate axes. However, under larger proper motion uncertainties, the influence of adopting incorrect Milky Way potential models diminishes, as the system deviates from its true configuration across all potential models.
@@
Thus, this study emphasizes the necessity of precise determination of proper motion, distances of satellite galaxies, and the Milky Way’s gravitational potential. We have seen how variations in these quantities can significantly shift our inference of the evolution of intrinsically stable satellite planes, especially in proper motion and halo mass of the host galaxy. We argue that underestimating the halo mass would also lead one to infer that more satellites escape from the host (and might thus appear unbound) than is the case in the true fiducial model.
@@
Accurate measurement of the Milky Way potential is thus vital for a comprehensive study of its satellite system and for broader insights into galactic evolution and interactions. We also found that the results are not very sensitive to the choice of the number of integrations; in our study, 600 backward integrations per model were sufficient.
@@
However, it is crucial to acknowledge certain limitations in this research. A significant caveat is the absence of consideration for the gravitational effects of massive satellite galaxies like the LMC. These celestial bodies, owing to their substantial mass, could affect the mass estimates of the Milky Way galaxy (Erkal et al., 2020), and exert significant gravitational forces that can profoundly influence the Milky Way and its surroundings (Garavito-Camargo et al., 2021; Pawlowski et al., 2022; Vasiliev, 2023), and thus can influence the stability of the satellite plane. The omission of these influences in the current study represents a notable limitation in our analysis. Additionally, this study assumes a spherical dark matter halo; however, it’s important to acknowledge that the halo could be triaxial (Law et al., 2009), which might have implications for our findings.
@@
Future studies could enhance our understanding by incorporating the gravitational influence of massive satellite galaxies like the LMC and SMC, as well as considering different shapes of the dark matter halo. Such an approach would allow for a more comprehensive analysis of the satellite plane’s dynamics, taking into account the complexities of real-world scenarios. Additionally, in this study, we assume a static dark matter halo over the 5 Gyr integration period. However, this simplification overlooks important evolutionary effects. Over time, the halo mass is expected to grow as it accretes dark matter, with a smaller mass in the past and a larger mass in the future. This impacts satellite orbits, particularly at larger radii. Pace et al. (2022) focuses on observations and finds that the orbital computations of 46 dwarf spheroidal galaxies (dSphs) using a static Milky Way potential reveal that, while pericenter and apocenter distances can be derived, the inclusion of the LMC’s potential alters these parameters by more than 25% for 40% of the sample. Similar studies have also been conducted using cosmological simulations D’Souza & Bell (2022); Santistevan et al. (2024). Additionally, the influence of the baryonic disk leads to adiabatic contraction, increasing the dark matter density at smaller radii, which affects the inner orbits of satellites (Cautun et al., 2020). Generally, however, such additional dynamical effects can be expected to increase the uncertainty in past satellite trajectories, as we have only limited constraints on their exact implementation. We therefore expect them to exacerbate the trend of inferring less stable satellite plane dynamics when backward-integrating. Moreover, an important dynamical process that is not included in our model is dynamical friction. This effect, especially for massive satellites orbiting within a host potential, can lead to the gradual decay of orbits as they lose energy and angular momentum to the dark matter background (Esquivel & Fuchs, 2007; Ogiya & Burkert, 2016; Taylor & Babul, 2001). Over time, this process can significantly alter satellite trajectories and contribute to the disruption of coherent planar structures.
@@
Our method can be extended to examine the observed planes of satellites of M31, Cen A and others. However, the application to these systems requires careful consideration of observational uncertainties, such as proper motions, distances, and gravitational potentials, which are significantly higher compared to our current analysis (in the case of M31) or thus far inaccessible (for more distant systems). Additionally, apart from these observational limitations, the evolution of such planes would also be sensitive to the initial distribution of the positions and velocities of the test satellites. Variations in the initial conditions could lead to notable differences in the structure and persistence of the planes over time. By accounting for these factors, our approach can provide valuable insights into the dynamics of satellite galaxies in diverse environments, further broadening the scope of our findings.
@@
Our numerical experiments demonstrate how intrinsically stable, co-orbiting planes of satellite galaxies display a significant increase in their plane height if they are backward-integrated after applying measurement errors. This is already the case for errors in proper motion comparable to the current Gaia systematics of $\sim 0.033,\mathrm{mas\,yr}^{-1}$ (Lindegren et al., 2021; McConnachie & Venn, 2020; Li et al., 2021). Such an increase in plane height was argued by Sawala et al. (2023) as indicative that the Milky Way’s Vast Polar Structure is more likely a transient rather than a rotationally supported feature and thus consistent with $\Lambda$CDM expectations. Similar arguments were put forward earlier by Maji et al. (2017) based on earlier, less accurate proper motion measurements. In their studies, the authors integrated the orbits of observed Milky Way satellite galaxies in a static potential, akin to our numerical experiments, and measured the time evolution of the inferred thickness of plane fits. However, our work shows that such an increasing thickness does not necessarily imply an absence of a kinematically coherent, dynamically stable, or long-lived satellite plane. In the case of an underlying, intrinsically stable satellite plane, realistic measurement errors alone, as well as assuming a mismatching Galactic potential, can already cause apparent thickening. Furthermore, one needs to be cautious in how errors are applied. Given that, for example, the measured most-likely proper motions of observed satellite galaxies are already subject to measurement error, sampling the errors in a Monte-Carlo fashion (as e.g. done in Sawala et al. 2023) then effectively applies these errors twice. For our study, this means that our applied proper motion error on the level of $\epsilon_{\mu}=0.04\,\mathrm{mas\,yr}^{-1}$ corresponds to a case of Monte-Carlo sampling of orbits based on observed data with a measurement error of $\epsilon_{\mu}/\sqrt{2}=0.028\,\mathrm{mas\,yr}^{-1}$. Monte-Carlo sampling thus further enhances the effect we have identified in our study of artificially increasing the inferred satellite plane height when backward-integrating, because it typically adds more dispersion to an ensemble of satellites. Merely finding an increase in the inferred plane height of an observed satellite structure under orbit integration – without accounting for the impact of measurement errors (especially in proper motions) or mismatches in the assumed Galactic gravitational potential – is thus insufficient to claim that the structure is not dynamically stable or transient, or that it is therefore consistent with $\Lambda$CDM.
@@ Appendix A Analysis at $\theta_{\max}=40^{\circ}$ and $60^{\circ}$
Fig. 15 and 11 display results analogous to those in Section 3. These scenarios exhibit qualitatively comparable patterns. The figures are segmented into four panels, each representing different levels of proper motion uncertainties. The mean $c/a$ ratios for forward integration are depicted with black curves, whereas those for backward integration are shown in colored curves. Similar to the previously discussed case, it becomes increasingly difficult to accurately determine the stability of the satellite plane as the level of proper motion errors escalates. Furthermore, in these results too, the behavior of the system under $MW_{\text{low}}$ is different from the other two potential models.
@@
Fig. 2 shows that as $\theta_{\text{max}}$ increases, the orbits become more eccentric, and this effect is clearly observed when we consider the results under $\theta_{\text{max}}$. By comparing the results at $\theta_{\text{max}}=40\%,60\%,80\%$, as shown in Tables 5, 6, and 4, we see that at a given level of proper motion uncertainty, the satellite plane widens more for larger $\theta_{\text{max}}$. For instance, in Model BI-02, when $\epsilon_{\mu}=0.04$ mas yr-1, $\Delta_{c/a}$ is 0.04, 0.06, and 0.07 for $\theta_{\text{tan}}=40\%,60\%,80\%$, respectively, and increases gradually. Furthermore, the influence of eccentric orbits is also apparent when comparing the curves in Fig. 15, 11, and 6. In cases with less eccentric orbits ($\theta_{\text{tan}}=40\%$), the plane appears relatively smoother and flatter. In contrast, as the orbits become more eccentric, the evolution of the plane flattening becomes increasingly irregular and pronounced, indicating a significant impact of orbital eccentricity on the satellite’s dynamics. This suggests that as $\theta_{\text{max}}$ increases, not only does the eccentricity of the orbits intensify, but the overall structure and stability of the orbital plane are also affected, leading to a more complex and varied orbital configuration.
@@
Similarly, Fig. 10 and 12 display the effects of distance uncertainties on the test satellite system for $\theta_{tan}$= 40°and 60°, respectively. The analysis, as discussed in Section 3, indicates that the inclusion of distance errors has a noticeable effect on the $c/a$. However, this does not significantly impact the overall evolution of $c/a$.
@@
Finally, similar to Fig. 9, here Fig. 13 and 14 display the number of test satellites that exceeded the 300 kpc radius threshold under $\theta_{max}$ = 40°and 60°, respectively. The effect of proper motion errors is depicted in each of the four panels, with each line within a panel representing the influence of different Milky Way potential models. Dotted-dashed curves show the effect of 5% distance uncertainties, while solid curves show 0% distance uncertainties. The influence of eccentricities is clearly apparent in these figures. For a given model, less eccentric orbits result in fewer test satellites escaping the system compared to those with higher eccentricities. This trend is further corroborated by the data presented in Tables 4, 5, and 6. In these tables, under all models, the column representing the number of satellites with distances greater than 300 kpc ($N_{sat}$ $>$ 300 kpc) consistently increases as the eccentricity of the orbits increases.
@@
Figure 10: Analogous to Fig 7, results for $\theta_{tan}$= 40°: Evolution of the plane of satellite galaxies with proper motion uncertainties included in each panel. The black dashed-dotted curves represent the mean axis ratio for forward integration, while the colored dashed curves indicate the mean axis ratio for backward integration under different potential models characterized by halo mass. Specifically, the blue, green, and red dashed curves correspond to the ${c/a}$ values under the $MW_{\text{low}}$, $MW_{\text{fiducial}}$, and $MW_{\text{high}}$ potential models, respectively. Note that in these results we account $5\%$ of distance uncertainties.
@@ B.1 Correlated PM and Dist errors
We also test whether a correlation between proper motion and distance errors affects the inferred evolution of the satellite plane, as fainter test satellites could have both larger proper motion and distance uncertainties. To investigate this, we performed an additional analysis, comparing two cases (see Fig. 16): one with a proper motion (PM) error of 0.04 mas/yr and a 10% distance error, and another with a PM error of 0.1 mas/yr and a 10% distance error. Each plot includes two backward integrations—one with a PM-distance error correlation and one without. In both cases, the same errors are used, but in the ’correlated’ case, the errors are ranked and assigned to the test satellites such that satellites farther away have progressively larger distance and proper motion errors. We find that the effect is not extreme but leads to an increase in the inferred widening of the satellite plane. This suggests that the approach of using uncorrelated errors is conservative and at most underestimates the impact of uncertainties. Moreover, as expected, with higher PM error, the effect becomes more pronounced, aligning with what we observed in our main results.

### v2

@@ 4 Conclusion and outlook
We investigated the impact of different types of measurement uncertainties on the inferred stability of artificial, intrinsically stable planes of satellite galaxies. Our key results are as follows.
@@
  • At the level of minimum Gaia uncertainties (with $\epsilon_{\mu}$ = 0.04 mas yr-1, similar to Gaia systematics), the inferred plane widths begin to diverge from the true value toward larger values. Although the thin initial plane remains relatively well constrained, its inferred width relative to the true plane almost doubles within 3 Gyr. As we increase the applied proper motion uncertainties ( $\epsilon_{\mu}$ = 0.08 and 0.12 mas yr-1), the inferred plane widths increase substantially, and one might infer that the intrinsically stable satellite planes are highly unstable. The $c/a$ shows a linear increase with the addition of more uncertainties.
@@
  • When performing backward integration with three different intrinsic plane heights ($h=10,20$, and 30 kpc), we find that applying realistic uncertainties in the proper motion introduces substantial broadening in the plane of the satellites in all cases. The apparent broadening effect we identified is therefore not sensitive to the exact intrinsic plane height assumed for the plane.
@@
  • When applying 5% distance uncertainties, we observe a noticeable effect that increases the plane width during backward integration relative to the intrinsic plane width. However, this effect does not dominate the inferred evolution of the satellite plane. Additionally, the impact of distance uncertainties becomes less important with higher levels of proper motion uncertainties.
@@
  • The adoption of incorrect Milky Way potential models leads to an apparent widening of the inferred past satellite planes. This impact is especially significant for $MW_{\text{low}}$, which has a 40% smaller halo mass compared to the fiducial model considered for our intrinsic satellite planes. Due to its lower mass, $MW_{\text{low}}$ struggles to maintain cohesion among the test satellites, leading to increased radial distances and an extension of its major, minor, and intermediate axes. However, under larger proper motion uncertainties, the influence of adopting incorrect Milky Way potential models diminishes, as the system deviates from its true configuration across all potential models.
@@
Thus, this study emphasizes the necessity of precise determination of proper motion, distances of satellite galaxies, and the Milky Way’s gravitational potential. Variations in these quantities can significantly shift our inference of the evolution of intrinsically stable satellite planes, especially with respect to proper motion and halo mass of the host galaxy. We argue that underestimating the halo mass also leads to the inference that more satellites escape from the host (and might thus appear unbound) than in the true fiducial model.
@@
Accurate measurement of the Milky Way potential is therefore vital for a comprehensive study of its satellite system and for broader insights into galactic evolution and interactions. We also find that the results are not very sensitive to the number of integrations; in our study, 600 backward integrations per model are sufficient.
@@
However, it is crucial to acknowledge certain limitations in this research. A significant caveat is the omission of the gravitational effects of massive satellite galaxies such as the LMC. These celestial bodies, owing to their substantial mass, can affect the mass estimates of the Milky Way galaxy (Erkal et al., 2020) and exert significant gravitational forces that profoundly influence the Milky Way and its surroundings (Garavito-Camargo et al., 2021; Pawlowski et al., 2022; Vasiliev, 2023), and thus the stability of the satellite plane. This omission represents a notable limitation in our analysis. Additionally, this study assumes a spherical dark matter halo; however, it is important to acknowledge that the halo could be triaxial (Law et al., 2009), which might have implications for our findings.
@@
Future studies could enhance our understanding by incorporating the gravitational influence of massive satellite galaxies, such as the LMC and SMC, as well as by considering different shapes of the dark matter halo. Such an approach would allow for a more comprehensive analysis of the satellite plane dynamics, taking into account the complexities of real-world scenarios. Additionally, in this study, we assume a static dark matter halo over the 5 Gyr integration period. However, this simplification overlooks important evolutionary effects. Over time, the halo mass is expected to grow as it accretes dark matter, resulting in a smaller mass in the past and a larger mass in the future. This affects satellite orbits, particularly at larger radii. Pace et al. (2022) focuses on observations and finds that the orbital computations of 46 dwarf spheroidal galaxies (dSphs) using a static Milky Way potential reveal that, while pericenter and apocenter distances can be derived, the inclusion of the LMC’s potential alters these parameters by more than 25% for 40% of the sample. Similar studies have also been conducted using cosmological simulations D’Souza & Bell (2022); Santistevan et al. (2024). Additionally, the influence of the baryonic disk leads to adiabatic contraction, increasing the dark matter density at smaller radii, which affects the inner orbits of satellites (Cautun et al., 2020). However, in general, such additional dynamical effects are expected to increase the uncertainty in past satellite trajectories, as we have only limited constraints on their exact implementation. We therefore expect them to exacerbate the trend of inferring less stable satellite plane dynamics when backward integrating. Moreover, our models does not include an important dynamical process: dynamical friction. This effect, especially for massive satellites orbiting within a host potential, can lead to the gradual decay of orbits as they lose energy and angular momentum to the dark matter background (Esquivel & Fuchs, 2007; Ogiya & Burkert, 2016; Taylor & Babul, 2001). Over time, this process can significantly alter satellite trajectories and contribute to the disruption of coherent planar structures.
@@
Our method can be extended to examine the observed planes of satellites of M31, Cen A, and others. However, the application to these systems requires careful consideration of observational uncertainties, such as proper motions, distances, and gravitational potentials, which are significantly greater compared to our current analysis (in the case of M31) or thus far inaccessible (for more distant systems). In addition to these observational limitations, the evolution of such planes is also sensitive to the initial distribution of the positions and velocities of the test satellites. Variations in the initial conditions could lead to notable differences in the structure and persistence of the planes over time. By accounting for these factors, our approach can provide valuable insights into the dynamics of satellite galaxies in diverse environments, further broadening the scope of our findings.
@@
Our numerical experiments demonstrate how intrinsically stable, co-orbiting planes of satellite galaxies display a significant increase in their plane height when backward integrated after applying measurement errors. This effect is already evident for proper motion errors comparable to the current Gaia systematics of $\sim 0.033,\mathrm{mas\,yr}^{-1}$ (Lindegren et al., 2021; McConnachie & Venn, 2020; Li et al., 2021). Sawala et al. (2023) argued that such an increase in plane height indicates that the Milky Way’s Vast Polar Structure is more likely a transient rather than a rotationally supported feature, and thus consistent with $\Lambda$CDM expectations. Similar arguments were previously put forward by Maji et al. (2017), based on less accurate proper motion measurements. In their studies, the authors integrated the orbits of observed Milky Way satellite galaxies in a static potential, akin to our numerical experiments, and measured the time evolution of the inferred thickness of plane fits. However, our work shows that such an increasing thickness does not necessarily imply an absence of a kinematically coherent, dynamically stable, or long-lived satellite plane. In the case of an underlying, intrinsically stable satellite plane, realistic measurement errors alone, as well as assuming a mismatching Galactic potential, can already cause apparent thickening. Furthermore, care must be taken in how errors are applied. For example, the measured most-likely proper motions of observed satellite galaxies are already subject to measurement error. Sampling the errors in a Monte Carlo fashion (e.g., as demonstrated by Sawala et al. 2023) then effectively applies these errors twice. In our study, this means that our applied proper motion error of $\epsilon_{\mu}=0.04\,\mathrm{mas\,yr}^{-1}$ corresponds to a Monte Carlo sampling of orbits based on observed data with a measurement error of $\epsilon_{\mu}/\sqrt{2}=0.028\,\mathrm{mas\,yr}^{-1}$. Monte Carlo sampling therefore enhances the effect we identified in our study of artificially increasing the inferred satellite plane height when backward integrating, because it typically adds more dispersion to an ensemble of satellites. Merely finding an increase in the inferred plane height of an observed satellite structure under orbit integration–without accounting for the impact of measurement errors (especially in proper motions) or mismatches in the assumed Galactic gravitational potential–is therefore insufficient to claim that the structure is not dynamically stable or transient, or that it is therefore consistent with $\Lambda$CDM.
@@ Appendix A Analysis at $\theta_{\max}=40^{\circ}$ and $60^{\circ}$
Figure 10 and 11 show results analogous to those in Section 3. These scenarios exhibit qualitatively comparable patterns. The figures are segmented into four panels, each representing different levels of proper motion uncertainties. The mean $c/a$ ratios for forward integration are depicted with black curves, whereas those for backward integration are shown in colored curves. Similar to the previously discussed case, it becomes increasingly difficult to accurately determine the stability of the satellite plane as the level of proper motion errors escalates. Furthermore, in these results too, the behavior of the system under $MW_{\text{low}}$ is different from the other two potential models.
@@
Figure 2 shows that as $\theta_{\text{max}}$ increases, the orbits become more eccentric, and this effect is clearly observed when we consider the results under $\theta_{\text{max}}$. By comparing the results at $\theta_{\text{max}}=40\%,60\%,80\%$, as shown in Tables 5, 6, and 4, we see that at a given level of proper motion uncertainty, the satellite plane widens more for larger $\theta_{\text{max}}$. For instance, in Model BI-02, when $\epsilon_{\mu}=0.04$ mas yr-1, $\Delta_{c/a}$ is 0.04, 0.06, and 0.07 for $\theta_{\text{tan}}=40\%,60\%,80\%$, respectively, and increases gradually. Furthermore, the influence of eccentric orbits is also apparent when comparing the curves in Fig. 10, 11, and 6. In cases with less eccentric orbits ($\theta_{\text{tan}}=40\%$), the plane appears relatively smoother and flatter. In contrast, as the orbits become more eccentric, the evolution of the plane flattening becomes increasingly irregular and pronounced, indicating a significant impact of orbital eccentricity on the satellite’s dynamics. This suggests that as $\theta_{\text{max}}$ increases, not only does the eccentricity of the orbits intensify, but the overall structure and stability of the orbital plane are also affected, leading to a more complex and varied orbital configuration.
@@
Similarly, Fig. 12 and 13 display the effects of distance uncertainties on the test satellite system for $\theta_{tan}$= 40°and 60°, respectively. The analysis, as discussed in Section 3, indicates that the inclusion of distance errors has a noticeable effect on the $c/a$. However, this does not significantly impact the overall evolution of $c/a$.
@@
Finally, similar to Fig. 9, here Fig. 14 and 15 display the number of test satellites that exceeded the 300 kpc radius threshold under $\theta_{max}$ = 40°and 60°, respectively. The effect of proper motion errors is depicted in each of the four panels, with each line within a panel representing the influence of different Milky Way potential models. Dash-dotted curves show the effect of 5% distance uncertainties, while solid curves show 0% distance uncertainties. The influence of eccentricities is clearly apparent in these figures. For a given model, less eccentric orbits result in fewer test satellites escaping the system compared to those with higher eccentricities. This trend is further corroborated by the data presented in Tables 4, 5, and 6. In these tables, under all models, the column representing the number of satellites with distances greater than 300 kpc ($N_{sat}$ $>$ 300 kpc) consistently increases as the eccentricity of the orbits increases.
@@
Figure 12: Analogous to Fig 7, results for $\theta_{tan}$= 40°: Evolution of the plane of satellite galaxies with proper motion uncertainties included in each panel. The black dashed-dotted curves represent the mean axis ratio for forward integration, while the colored dashed curves show the mean axis ratio for backward integration under different potential models characterized by halo mass. Specifically, the blue, green, and red dashed curves correspond to the ${c/a}$ values under the $MW_{\text{low}}$, $MW_{\text{fiducial}}$, and $MW_{\text{high}}$ potential models, respectively. Note that in these results we account $5\%$ of distance uncertainties.
@@ B.1 Correlated PM and Dist errors
We also test whether a correlation between proper motion and distance errors affects the inferred evolution of the satellite plane, as fainter test satellites could have both larger proper motion and distance uncertainties. To investigate this, we performed an additional analysis, comparing two cases (see Fig. 16): one with a proper motion error of 0.04 mas yr-1 and a 10% distance error, and another with a PM error of 0.1 mas yr-1 and a 10% distance error. Each plot includes two backward integrations—one with a PM-distance error correlation and one without. In both cases, the same errors are used, but in the ’correlated’ case, the errors are ranked and assigned to the test satellites such that satellites farther away have progressively larger distance and proper motion errors. We find that the effect is not extreme but leads to an increase in the inferred widening of the satellite plane. This suggests that the approach of using uncorrelated errors is conservative and at most underestimates the impact of uncertainties. Moreover, as expected, with higher PM error, the effect becomes more pronounced, aligning with what we observed in our main results.